OpenStudy (anonymous):

i will give a meadle and fan !!!!! A car travels 132 miles using 4 gallons of gas.AT that rate, how far can the car travel using 7 gallons of gas. A. 231 B.264 C.134 D.1,260

OpenStudy (anonymous):

A. 132/4=33 33*7=231

OpenStudy (anonymous):

THANS MUSIC LIFE

OpenStudy (anonymous):

You're welcome! :) @oregon234

OpenStudy (anonymous):

\(\Large \frac{4/4}{132/4}\) = \(\Large \frac{1}{33}\) <- That tells you he is going 33 miles per gallon, so if you multiply 33 by 7, that is your answer